OpenStudy (anonymous):

Choose the equation that could be used to find two consecutive integers whose sum is 67. n + (n + 2) = 67 n − (n + 1) = 67 n + n = 67 n + (n + 1) = 67

OpenStudy (anonymous):

define consecutive

OpenStudy (anonymous):

Numbers which follow each other in order, without gaps, from smallest to largest.

OpenStudy (anonymous):

define sum

OpenStudy (anonymous):

add

OpenStudy (anonymous):

ok so if n is the starting number the next consecutive number would be n+1, right?

OpenStudy (anonymous):

i believe so

OpenStudy (anonymous):

so which one do you think is the answer

OpenStudy (anonymous):

b.

OpenStudy (anonymous):

what does sum mean again?

OpenStudy (anonymous):

O sorry i mean D.

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ty :)

OpenStudy (anonymous):

It would be n+(n+1)=67 We know for 2 numbers to be consecutive, one has to be one more than the other (i.e. 1, and 2 are consecutive numbers, where 2 is 1 greater than 1) So, n+(n+1)=67 will be applied :) Now solve: 2n+1=67 2n=67-6 2n=66 n=66/2 n=33 So, the other number will be (x+1)=33+1=34 The first number is 33 and second is 34 I hope I'm clear :) if not then just lemme know I'll explain it again :)

OpenStudy (anonymous):

good job @xoya , however it never asked for n

OpenStudy (anonymous):

Thankyou :) I wanted to clear so I explained it :)

OpenStudy (anonymous):

Ty to Xoya

OpenStudy (anonymous):

Your welcome :)